OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1). − −

OpenStudy (anonymous):

7/3 or 2 and 1/3

OpenStudy (bahrom7893):

Slope: M=(y2-y1)/(x2-x1)=(1+6)/(-3+6)=7/3